# Electromagnetic energy calibration of the SoLid detector with horizontal muons

### v1

@@ 1 Introduction
The basic detection unit of the detector is a 5 cm side polyvinyl toluene (PVT) cube. PVT is a cheap plastic scintillator with linear response over a wide range of energy. Each cube has two neutron detection screens (microcomposite 6LiF:ZnS(Ag)) placed on adjacent faces. The detection units are combined into planes of 16$\times$16 units each. Each cube is individually wrapped in Tyvek to prevent scintillation light from escaping. Furthermore, each plane is surrounded by two square Tyvek sheets, which optically decouple the planes by further preventing the passage of light between them. The latter is a very important feature for both calibration and reconstruction, since it simplifies a problem from 3D to 2D. Ten planes make up a module, and there are five modules in total. The scintillation light from the PVT cube is collected by two vertical and two horizontal WLS fibres that pass through each detection cell. One side of each WLS fibre is covered with a Mylar foil that acts as a mirror to reflect the incoming light. The second side is coupled with multi-pixel photon counters (MPPCs) that read out the light. The digitised version of the readout is the initial input received from the detector. Figure 1 shows the schematic view of the SoLid detection unit and the detection plane. The origin of the coordinate system corresponds to the position of the reactor core. The $x-y$ coordinates are defined by the orientation of the detector planes, with $x$ being the horizontal direction and $y$ the vertical, respectively. The $z$ axis is perpendicular to the detector planes.
@@
However, the detector does not directly deliver the positron energy information at the cube level. The starting point for any analysis is the digitised readout from the MPPCs. Therefore, signal candidates must be defined from this input. More precisely, the MPPC response must be transformed back into the list of detection cubes involved in the event. This list of cubes contains the energy deposits from the annihilation gamma and positron if the mean path length of the annihilation gamma is enough to escape the annihilation cube. In addition, to verify the energy estimator suggested in Equation 1.1, the reconstructed annihilation gamma clusters also provide a very powerful background rejection tool. Figure 2 sketches the basics of the reconstruction problem. If there is an energy deposit $E_{j}$, then the WLS fibres act as linear projectors to transfer light to the MPPC. As mentioned above, the planes are optically decoupled; hence, the reconstruction problem can be posed for each plane individually. By postulating $a_{ij}$ as the projector from cube j to MPPC i, the readout value for this particular MPPC is calculated as follows:
@@
where $p$ is the column vector of the readout projections. $E$ is a column vector of unknown energy deposits that are determined by the reconstruction procedure. Finally, $A$ is a matrix of 64 $\times$ 256 dimensions. It is called the System Matrix and embodies the best of our knowledge about the detector behaviour at each stage, from the light generation to the digitisation. Finally, it must be complemented with the absolute energy scale to transform the energy of the cubes from the analogue-to-digital converter units (ADCs) to the physics units (MeV). The description of the methods for solving Equation 1.3 is beyond the scope of this article. The tools available on the market, together with the baseline choice made for the SoLid experiment, known as the CCube algorithm, are exhaustively discussed in  [7]. The equation in the paper is solved assuming that the System Matrix is determined from elsewhere. The derivation of this matrix and the determination of the absolute energy scale are the subjects of this article.
@@ 2 Cosmic muons as the calibration source
The projector values $a_{ij}$ in Equation 1.2 are not constrained a priori. Their initial features are obtained from a simplified optical simulation. The simulation is performed with the GODDeSS extension framework [8] for Geant4. It consists of one row of 16 identical PVT cubes without neutron detection screens, but wrapped with reflective Tyvek paper. Both horizontal and vertical WLS fibres are paired with MPPCs, which read out the number of arriving scintillation photons. A single muon is generated such that it crosses the central cube of the system. The setup is sketched in Figure 3.
@@
First, the simulation shows that the scintillation light is not contained within a single cube. A significant fraction of photons ($\sim$10%) is propagated to the neighbouring cubes, in particular through the holes created by the WLS positioning tolerances. This effect will be referred to as Light Leaks (LL) in the following. The fraction of photons that end up two cubes away from the crossed cube is at a subpercent level, and thus neglected. Therefore, each energy deposit has only 12 non-zero projector values associated with it. Four of them are related to the fibres in the main cube, where the energy deposit has occurred, and another eight are related to the four neighbouring cubes (2 on the horizontal axis and 2 on the vertical) created by the LL. Secondly, the simulation reveals asymmetries in light sharing between the fibres. These asymmetries are observed between the horizontal and vertical fibres. Moreover, the amount of light shared between the two horizontal (vertical) fibres that cross the same cube is not equivalent either. Thus, a per-fibre determination of the System Matrix elements, i.e. calibration, is desirable. The final stage of the simplified optical simulation scrutiny is related to the impact of the WLS groove size. The original size of the side of the groove is 5 mm, while the diameter of the fibre is 3 mm. In the modified version of the simulation, the side of the groove is changed to match the size of the WLS fibre, so that there is no air gap between the two. This modification demonstrated a drastic change in the distributions of scintillating light, which, in turn, indicates miscalibrations due to the actual geometry of the detector. This effect also varies from cube to cube. Other effects such as temperature or the detector displacements triggered by the calibration sources’ movements can also influence the detector response. The modelling of the miscalibrations cannot therefore rely solely on physics assumptions such as attenuation length or coupling effect. Since all detection units are identical in the Geant4 simulation, the modelling cannot rely on it either. There are two ways to address this problem: develop a much more detailed simulation or alternatively use data directly.
@@
where $p_{i}$ is the direct input of the detector. Hence, if the total energy deposited in the plane is known, the projector value $a_{ij}$ is the only unknown. Therefore, in such a case, all 12 projector values associated with the cube, where the energy deposit occurs, can be measured directly. The horizontal cosmic muons meet all the requirements listed above. On average, muons deposit 1.6 MeV/cm (which translates to 8 MeV for the SoLid detection unit size of 5 cm). As such, it is possible to accurately measure both the light-sharing properties within the main cube traversed by a muon and in the neighbouring cubes impacted by the LL. Furthermore, the muon track typically crosses several cubes in the detector. Thus, it is possible to use a single track to calibrate multiple cubes. However, to fit the condition of the single energy deposit in the plane, only a subset of the cosmic muon sample is considered for calibration purposes. It consists of parts of the muon tracks that are horizontal enough to cross only one cube in a plane.
@@ 3 Selection of Horizontal Muons
Horizontal muon candidates are selected from both reactor-off (ROff) and reactor-on (ROn) samples. The selection itself is implemented in the SoLid Analysis Framework (Saffron2). The main objective of the tool is to cluster the waveforms recorded by detectors that are close in time and space. The waveforms within a cluster are divided into three mutually exclusive categories: muons, nuclear signals, and electromagnetic signals. The latter two are aimed at selecting the signals originating from the IBD neutron (nuclear) and positron together with annihilation gammas (electromagnetic), respectively. The exhaustive description of the Saffron2 software can be found in [9]. As for muon clusters, they are searched for in the first place that satisfies a single requirement on the fibre read-out: the presence of high-energy channels with an amplitude greater than 200 ADC counts. If the number of these channels is less than 11, the muon candidate is tagged as a Type 0 muon. In most cases, Type 0 muons correspond to the so-called clipping muons, which cross only a few cubes on the edge of the detector. In other cases, the muon candidate deposits energy in a larger number of cubes that most likely form a track. The latter assumption is checked with two weighted least-squares linear fits (vertical and horizontal projections). The candidates are further divided into two categories according to the convergence of the fit: Type 1 if one of the fits has failed; Type 2 if both fits were successful. The horizontal muons used for the calibration are selected exclusively within the Type 2 muon sample. It is important to mention that the fits are performed solely with the fibre-level information, i.e. the reconstruction algorithm is not involved in determining the positions of the cubes on the track. Therefore, one can simultaneously solve equation 1.3 and determine the System Matrix. Several additional requirements are applied to Type 2 muons to increase the quality of the sample and ensure that there is indeed a single cube hit in the plane. First, the start and end cubes of the track have to be on the detector border. This requirement allows the muons, which are captured while crossing the detector (stopping muons) and as such have a different stopping power, to be rejected. Second, the muon track length must be greater than or equal to 7 cubes. This rejects candidates that are likely excited nuclei. Third, the slopes $x/z$ and $y/z$ must provide $\cos\theta$ less than 0.8, where $\theta$ is the polar angle of the muon track (approximately 40). The development of the muon selection is discussed in [10]. An example of a Type 2 muon track that meets all the listed conditions is shown in Figure 4. However, not all cubes along this track are used in the calibration procedure. To select the planes in which the muon has indeed deposited its energy in a single detection unit, the cube is kept if and only if the reconstructed track enters and exits the cube through a fiducial window. This window is shown in pink in Figure 5. The set of such cubes from the dedicated muon tracks is referred to as horizontal muons in the following.
@@ 5 Absolute calibration
The results of the relative calibration must be complemented by the absolute energy scale. This is a factor that converts the energy measurement from units provided by the electronics (ADC) to physics units (MeV). One possibility of determining it consists in using the energy-loss distribution from the cosmic horizontal muon. Since the relative calibration provides equalised responses from all the individual detection units, it is enough to compute a single conversion factor instead of proceeding cube-by-cube. All the energy loss values are therefore gathered in a single distribution, to which the Landau function convoluted with the Gaussian is fit. The most probable value of Landau gives the reference value in ADC units 111technically in pixel avalanches, or PA, but since the gain is equalised for all the fibres, the conversion factor from PA to ADC is known. The value in MeV is obtained from a dedicated Geant4 simulation. It includes the full-scale detector geometry that is combined with the CRY engine [14] to simulate muons. All energy loss values obtained from the simulation are also combined in a single distribution. The ratio between the most probable values extracted from two fits defines the absolute energy scale.
@@ 5.1 Absolute calibration with the americium-beryllium source
A novel method (to our knowledge) has been introduced by making use of an americium-beryllium (AmBe) radioactive calibration source that was used in SoLid to calibrate the neutron signal. The idea consists of measuring the response of the detector to the 4.44 MeV gamma emitted by excited 12C, with a branching fraction of 60%. These gammas can experience an electron-positron ($e^{+}e^{-}$) pair conversion that provides a well-defined monoenergetic calibration source of 3.4 MeV. Although Compton scattering is the dominant process, the conversion appears in 3.5% of the cases according to a generic Geant4 simulation (with the G4EmStandardPhysics physics list). The presence of the positron makes the calibration signature identical to that of the IBD signal candidates. Therefore, all tools developed for the topological identification of IBD events are highly relevant and can be used to select $e^{+}e^{-}$.
@@
$\begin{gathered}^{241}Am\;\longrightarrow\;\alpha\;+\;^{237}N\!p\\ \alpha\;+\;^{9}Be\;\longrightarrow\;n\;+\;^{12}C^{*}\\ {}^{12}C^{*}\;\longrightarrow\;^{12}C\;+\;\gamma\;(4.438\;\textnormal{MeV})\;.\end{gathered}$ (5.1)
Furthermore, neutrons issued from the capture of $\alpha$ in ${}^{9}Be$, can interact with 12C nuclei in the detector materials. These neutrons are energetic enough to excite the 12C nucleus. During further deexcitation, the same 4.44 MeV gamma can emerge as in the nominal decay chain. The initial sample of the calibration signal is therefore increased by 15%.
@@
The green signal distribution in Figure 12 shows a sharp peak at 3.4 MeV of energy deposited by the $e^{+}e^{-}$ conversion pair. A small shoulder on the right-hand side of this peak is observed and corresponds to additional energy deposits from the annihilation gammas experiencing Compton scattering in the AC. The topological selection of calibration candidates (as performed with IBD signal candidates discussed in [7] (two back-to-back photons separated by several cubes) removes most of this contribution.
This Geant4 simulation is processed with the SoLid reconstruction software Saffron2. Basic physics information about the event is built, such as the additional cubes in the envelope, the number of reconstructed annihilation gamma candidates, and the cosine of the angle between the annihilation gamma candidates  [^2] The processing is done with the calibration based on the use of the horizontal muons and within the special module called the ReadOut Simulation or ROSim. The goal of this module is to make the Geant4 output files compatible with the Saffron2 input format. ROSim transforms the Geant4 energy deposits into the number of scintillation photons that are read out by the correspondent MPPCs. However, it skips the actual optical simulation and generation of the waveforms to optimise the computational cost of the processing. The detailed tuning of the ROSim is reported in [15].. Both signal and background events are preselected according to the geometric requirements : AC must be a unique cube in the envelope, the number of reconstructed annihilation gamma candidates must be 2, and they must be located in different hemispheres of the detector.
@@ 5.1.1 Calibration method
Figure 15: An example of the unbinned event-by-event ML fit performed on the AmBe ROSim sample. The background component is displayed in red, the signal component in blue, the total model in black, and the data are represneted by the black points.
@@ 5.1.2 Uncertainty budget
In order to evaluate how robust the determination of the absolute energy scale factor is to an additional background, the scan of the best pair of parameters is performed in the presence of a linearly-decreasing background component with different global contamination fractions (5, 10 or 20%). The results are collected in Table 1. The determination of the energy scale factor is stable even with the highest contamination, suggesting that the systematic budget can be close to the statistical uncertainty. In contrast, the energy resolution parameter strongly depends on the presence of this additional potential background. This cannot thus be used reliably to determine the intrinsic energy resolution of the SoLid detector.

### v2

@@ 1 Introduction
The basic detection unit of the detector is a 5 cm side polyvinyl toluene (PVT) cube. PVT is a cheap plastic scintillator with a linear response over a wide range of energy. Each cube has two neutron detection screens (microcomposite 6LiF:ZnS(Ag)) placed on adjacent faces. The detection units are combined into planes of 16$\times$16 units each. Each cube is individually wrapped in Tyvek to prevent scintillation light from escaping. Furthermore, each plane is surrounded by two square Tyvek sheets, which optically decouple the planes by further preventing the passage of light between them. The latter is a very important feature for both calibration and reconstruction, since it simplifies the problem from 3D to 2D. Ten planes make up a module, and the detector comprises five modules in total. The scintillation light from the PVT cube is collected by two vertical and two horizontal WLS fibres that pass through each detection cell. One side of each WLS fibre is covered with a Mylar foil that acts as a mirror to reflect the incoming light. The second side is coupled with multi-pixel photon counters (MPPCs) that read out the light. The digitised version of the readout is the initial input received from the detector. Figure 1 shows the schematic view of the SoLid detection unit and the detection plane. The origin of the coordinate system corresponds to the position of the reactor core. The $x-y$ coordinates are defined by the orientation of the detector planes, with $x$ being the horizontal direction and $y$ the vertical, respectively. The $z$ axis is perpendicular to the detector planes.
@@
However, the detector does not directly deliver the positron energy information at the cube level. The starting point for any analysis is the digitised readout from the MPPCs. Therefore, signal candidates must be defined from this input. More precisely, the MPPC response must be transformed back into the list of detection cubes involved in the event. This list of cubes contains the energy deposits from the annihilation gamma and positron if the mean path length of the annihilation gamma is enough to escape the annihilation cube. In addition to verifying the energy estimator suggested in Equation 1.1, the reconstructed annihilation gamma clusters also provide a very powerful background rejection tool. Figure 2 sketches the basics of the reconstruction problem. If there is an energy deposit $E_{j}$, then the WLS fibres act as linear projectors to transfer light to the MPPC. As mentioned above, the planes are optically decoupled; hence, the reconstruction problem can be posed for each plane individually. By postulating $a_{ij}$ as the projector from cube j to MPPC i, the readout value for this particular MPPC is calculated as follows:
@@
where $p$ is the column vector of the readout projections. $E$ is a column vector of unknown energy deposits that are determined by the reconstruction procedure. Finally, $A$ is a matrix of 64 $\times$ 256 dimensions. It is called the System Matrix and embodies the best of our knowledge about the detector behaviour at each stage, from the light generation to the digitisation. Finally, it must be complemented with the absolute energy scale to transform the energy of the cubes from the analogue-to-digital converter units (ADCs) to the physics units (MeV). The description of the methods for solving Equation 1.3 is beyond the scope of this article. The tools available on the market, together with the baseline choice made for the SoLid experiment, known as the CCube algorithm, are exhaustively discussed in  [7]. The equation is solved assuming that the System Matrix is determined from elsewhere. The derivation of this matrix and the determination of the absolute energy scale are the subjects of this article.
@@ 2 Cosmic muons as the calibration source
The projector values $a_{ij}$ in Equation 1.2 are not constrained a priori. Their initial features are obtained from a simplified optical simulation. The simulation is performed with the GODDeSS extension framework [8] for Geant4. It consists of one row of 16 identical PVT cubes without neutron detection screens, but wrapped in reflective Tyvek paper. Both horizontal and vertical WLS fibres are paired with MPPCs, which read out the number of arriving scintillation photons. A single muon is generated such that it crosses the central cube of the system. The setup is sketched in Figure 3.
@@
First, the simulation shows that the scintillation light is not contained within a single cube. A significant fraction of photons ($\sim$10%) is propagated to the neighbouring cubes, in particular through the holes created by the WLS positioning tolerances. This effect will be referred to as Light Leaks (LL) in the following. The fraction of photons that end up two cubes away from the crossed cube is at a subpercent level, and thus neglected. Therefore, each energy deposit has only 12 non-zero projector values associated with it. Four of them are related to the fibres in the main cube, where the energy deposit has occurred, and another eight are related to the four neighbouring cubes (2 on the horizontal axis and 2 on the vertical) created by the LL. Secondly, the simulation reveals asymmetries in light sharing between the fibres. These asymmetries are observed between the horizontal and vertical fibres. Moreover, the amount of light shared between the two horizontal (vertical) fibres that cross the same cube is not equivalent either. Thus, a per-fibre determination of the System Matrix elements, i.e. calibration, is desirable. The final stage of the simplified optical simulation scrutiny is related to the impact of the WLS groove size. The width of the groove is 5 mm, while the diameter of the fibre is 3 mm. In the modified version of the simulation, the width of the groove is changed to match the size of the WLS fibre, so that there is no air gap between the two. This modification demonstrated a drastic change in the distributions of scintillating light, which, in turn, indicates miscalibrations due to the actual geometry of the detector. This effect also varies from cube to cube. Other effects such as temperature or the detector displacements triggered by the movement of the calibration sources can also influence the detector response. The modelling of the miscalibrations cannot therefore rely solely on physics assumptions such as attenuation length or coupling effect. Since all detection units are identical in the Geant4 simulation, the modelling cannot rely on it either. There are two ways to address this problem: develop a much more detailed simulation or alternatively use data directly.
@@
where $p_{i}$ is the direct input of the detector. Hence, if the total energy deposited in the plane is known, the projector value $a_{ij}$ is the only unknown. Therefore, in such a case, all 12 projector values associated with the cube, where the energy deposit occurs, can be measured directly. Cosmic muons can meet all the requirements listed above. On average, muons deposit 1.6 MeV/cm (which translates to 8 MeV for the SoLid detection unit size of 5 cm). As such, it is possible to accurately measure both the light-sharing properties within the main cube traversed by a muon and in the neighbouring cubes impacted by the LL. Furthermore, the muon track typically crosses several cubes in the detector. Thus, it is possible to use a single track to calibrate multiple cubes. However, to fit the condition of the single energy deposit in the plane, only a subset of the cosmic muon sample is considered for calibration purposes. It consists of parts of the muon tracks that are horizontal enough to cross only one cube in a plane.
@@ 3 Selection of Horizontal Muons
Horizontal muon candidates are selected from both reactor-off (ROff) and reactor-on (ROn) samples. The selection itself is implemented in the SoLid Analysis Framework (Saffron2). The main objective of the tool is to cluster the waveforms recorded by detectors that are close in time and space. The waveforms within a cluster are divided into three mutually exclusive categories: muons, nuclear signals, and electromagnetic signals. The latter two are aimed at selecting the signals originating from the IBD neutron (nuclear) and positron together with annihilation gammas (electromagnetic), respectively. The exhaustive description of the Saffron2 software can be found in [9]. As for muon clusters, they are searched for in the first place that satisfies a single requirement on the fibre read-out: the presence of high-energy channels with an amplitude greater than 200 ADC counts. If the number of these channels is less than 11, the muon candidate is tagged as a Type 0 muon. In most cases, Type 0 muons correspond to the so-called clipping muons, which cross only a few cubes on the edge of the detector. In other cases, the muon candidate deposits energy in a larger number of cubes that most likely form a track. The latter assumption is checked with two weighted least-squares linear fits (vertical and horizontal projections). The candidates are further divided into two categories according to the convergence of the fits: Type 1 if either one of the fits has failed; Type 2 if both fits were successful. The horizontal muons used for the calibration are selected exclusively from the Type 2 muon sample. It is important to mention that the fits are performed solely with the fibre-level information, i.e. the reconstruction algorithm is not involved in determining the positions of the cubes on the track. Therefore, one can simultaneously solve equation 1.3 and determine the System Matrix. Several additional requirements are applied to Type 2 muons to increase the quality of the sample and ensure that there is indeed a single cube hit in the plane. First, the start and end cubes of the track have to be on the detector border. This requirement allows the muons, which are captured while crossing the detector (stopping muons) and as such have a different stopping power, to be rejected. Second, the muon track length must be greater than or equal to 7 cubes. This rejects candidates that are likely excited nuclei. Third, the slopes $x/z$ and $y/z$ must provide $\cos\theta$ less than 0.8, where $\theta$ is the polar angle of the muon track (approximately 40). The development of the muon selection is discussed in [10]. An example of a Type 2 muon track that meets all the listed conditions is shown in Figure 4. However, not all cubes along this track are used in the calibration procedure. To select the planes in which the muon has indeed deposited its energy in a single detection unit, the cube is kept if and only if the reconstructed track enters and exits the cube through a fiducial window. This window is shown in pink in Figure 5. The set of such cubes from the dedicated muon tracks is referred to as horizontal muons in the following.
@@ 5 Absolute calibration
The results of the relative calibration must be complemented by the absolute energy scale. This is a factor that converts the energy measurement from units provided by the electronics (ADC) to physics units (MeV). One possibility of determining it consists of using the energy-loss distribution from the cosmic horizontal muon. Since the relative calibration provides equalised responses from all the individual detection units, it is sufficient to compute a single conversion factor instead of proceeding cube-by-cube. All the energy loss values are therefore gathered in a single distribution, to which the Landau function convoluted with the Gaussian is fit. The most probable value of Landau gives the reference value in ADC units 111technically in pixel avalanches, or PA, but since the gain is equalised for all the fibres, the conversion factor from PA to ADC is known. The value in MeV is obtained from a dedicated Geant4 simulation. It includes the full-scale detector geometry that is combined with the CRY engine [14] to simulate muons. All energy loss values obtained from the simulation are also combined in a single distribution. The ratio between the most probable values extracted from two fits defines the absolute energy scale.
@@ 5.1 Absolute calibration with the americium-beryllium source
A novel method (to our knowledge) has been introduced by making use of an americium-beryllium (AmBe) radioactive calibration source that was used in SoLid to calibrate the neutron signal. The idea consists of measuring the response of the detector to the 4.44 MeV gamma emitted by excited 12C, with a branching fraction of 60%. These gammas can experience an electron-positron ($e^{+}e^{-}$) pair conversion which provides a well-defined monoenergetic calibration source of 3.4 MeV. Although Compton scattering is the dominant process, the conversion appears in 3.5% of the cases according to a generic Geant4 simulation (with the G4EmStandardPhysics physics list). The presence of the positron makes the calibration signature identical to that of the IBD signal candidates. Therefore, all tools developed for the topological identification of IBD events are highly relevant and can be used to select $e^{+}e^{-}$.
@@
[…]
Furthermore, neutrons issued from the capture of $\alpha$ in 9Be, can interact with 12C nuclei in the detector materials. These neutrons are energetic enough to excite the 12C nucleus. During further deexcitation, the same 4.44 MeV gamma can emerge as in the nominal decay chain. The initial sample of the calibration signal is therefore increased by 15%.
@@
The green signal distribution in Figure 12 shows a sharp peak at 3.4 MeV of energy deposited by the $e^{+}e^{-}$ conversion pair. A small shoulder on the right-hand side of this peak is observed and corresponds to additional energy deposits from the annihilation gammas experiencing Compton scattering in the AC. The topological selection of calibration candidates of two back-to-back photons separated by several cubes (as performed with IBD signal candidates discussed in [7]) removes most of this contribution.
This Geant4 simulation is processed with the SoLid reconstruction software Saffron2. Basic physics information about the event, such as the additional cubes in the envelope, the number of reconstructed annihilation gamma candidates, and the cosine of the angle between the annihilation gamma candidates, is built  [^2] The processing is done with the calibration based on the use of the horizontal muons and within the special module called the ReadOut Simulation or ROSim. The goal of this module is to make the Geant4 output files compatible with the Saffron2 input format. ROSim transforms the Geant4 energy deposits into the number of scintillation photons that are read out by the correspondent MPPCs. However, it skips the actual optical simulation and generation of the waveforms to optimise the computational cost of the processing. The detailed tuning of the ROSim is reported in [15].. Both signal and background events are preselected according to the geometric requirements : AC must be a unique cube in the envelope, the number of reconstructed annihilation gamma candidates must be 2, and they must be located in different hemispheres of the detector.
@@ 5.1.1 Calibration method
Figure 15: An example of the unbinned event-by-event ML fit performed on the AmBe ROSim sample. The background component is displayed in red, the signal component in blue, the total model in black, and the data are represented by the black points.
@@ 5.1.2 Uncertainty budget
In order to evaluate how robust the determination of the absolute energy scale factor is to an additional background, the scan of the best pair of parameters is performed in the presence of a linearly-decreasing background component with different global contamination fractions (5, 10 or 20%). The results are collected in Table 1. The determination of the energy scale factor is stable even with the highest contamination, suggesting that the systematic budget is comparable to the statistical uncertainty. In contrast, the energy resolution parameter strongly depends on the presence of this additional potential background. This cannot thus be used reliably to determine the intrinsic energy resolution of the SoLid detector.